### v1

# Rule-based Generation of de Bruijn Sequences: Memory and Learning

Abstract
We investigate binary sequences generated by non-Markovian rules with memory length $\mu$, similar to those adopted in Elementary Cellular Automata. This generation procedure is equivalente to a shift register and certain rules produce sequences with maximal periods, known as de Bruijn sequences. We introduce a novel methodology for generating de Bruijn sequences that combines: (i) a set of derived properties that significantly reduce the space of feasible generating rules, and (ii) a neural network-based classifier that identifies which rules produce de Bruijn sequences. Experiments for large values of $\mu$ demonstrate the approach’s effectiveness and computational efficiency.
1 Introduction
Cellular Automata (CA) are a class of dynamical systems that evolve in discrete time and space (Wolfram1983, ; Wolfram1984, ). A CA consists of cells (or agents), each of which can adopt a state from a discrete set. The most commonly used alphabet is binary, typically denoted by $\{0,1\}$. The CA evolves by applying generating rules at each time step. These rules can be static or dynamic (i.e., they may change depending on the current state of the entire CA), and either local or global, depending on whether they involve only neighboring cells or the entire population.
From an initial configuration, the state of the CA evolves over time according to the specified rules. The update rule for each cell can depend on both the past states of the cell itself (temporal dimension) and the states of its neighboring cells (spatial dimension). Elementary one-dimensional cellular automata (1dCA) typically assume a memoryless structure with a linear topology, where each cell interacts only with its nearest neighbors (Wolfram1983, ). More complex models incorporate memory—i.e., dependence on a set of past states—and long-range spatial influences (Li, ; Ramon, ).
Relatively less attention has been given to zero-dimensional cellular automata (0dCA), which consist of a single cell whose state is updated based solely on its own past states. Despite their apparent simplicity, 0dCA do not involve interactions between multiple cells, their time evolution yields interesting sequences of symbolic states. The rules governing such systems depend on the memory $\mu$—the number of previous states influencing the next state—and the alphabet $\{\alpha\}$, the number of possible states a cell can take. Generally, the complexity of the system increases with both $\mu$ and $\alpha$. For simplicity, we focus on binary systems, i.e., $\alpha=2$, using the symbols $\{0,1\}$.
2 0-dimensional CA
Symbolic sequences naturally arise in discrete dynamical systems, e.g. CA (Jin, ). From an initial condition, the next state is computed according to a generation rule (or function) that depends on previous states. For instance, a classical non-Markovian binary time series can be defined by initial values $a_{1}$ and $a_{2}$, and a recursive rule:
$a_{k+2}=(a_{k+1}+a_{k})\mod 2,\quad\text{for }k=1,2,\ldots$ (1)
With initial states $a_{1}=0$ and $a_{2}=1$, the sequence continues as $a_{3}=1$, $a_{4}=0$, $a_{5}=1$, $a_{6}=1$, etc., producing the binary sequence:
[…]
This sequence is clearly periodic with period $T_{1}=3$. In this case, the next digit depends on the two previous digits, implying a memory $\mu=2$.
Other examples with $\mu=2$ can be constructed using logical operators. For instance, applying the AND function with the rules:
1 11,1 00,0 10,0 00formulae-sequence111formulae-sequence100formulae-sequence0100001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 0$\quad$ 0\,1$\rightarrow$ 0$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 0 , 0 1 → 0 , 0 0 → 0
to the initial pair 0 1010\,10 1 produces the sequence:
[…]
Similarly, applying the OR function:
1 11,1 01,0 11,0 00formulae-sequence111formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
to the same initial configuration 0 1010\,10 1 results in:
[…]
Both sequences converge to fixed values (0 or 1), i.e., their periods are $T_{2}=T_{3}=1$, and these fixed points are independent of the initial configuration.
In contrast, applying the XOR function:
1 10,1 01,0 11,0 00formulae-sequence110formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 0$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 0 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
to the initial state $a_{1}=0,a_{2}=1$ yields:
[…]
with period $T_{4}=3$. However, using the same rule with the initial configuration: $a_{1}=0,a_{2}=0$ results in a constant sequence $a_{k}=0$ for all $k>2$. These examples illustrate that the resulting behavior depends not only on the rule but also on the initial conditions.
For any finite memory $\mu<\infty$, all sequences generated in this framework are eventually periodic. That is, for any rule and initial condition, the sequence will eventually settle into a repeating pattern (also referred to as a motif or loop). The maximum possible period for a sequence with memory $\mu$ is:
$T_{\text{max}}=2^{\mu}$ (2)
As $\mu$ increases, the sequences may resemble aperiodic ones, making them useful for computing pseudorandom numbers to be applied in Monte Carlo simulations, and cryptographic systems (Etzion, ). Truly aperiodic behavior can only arise when $\mu\rightarrow\infty$, or when the alphabet is infinite (e.g., using real-valued states).
To explore periodicity in 0dCA, we adopt combinatorial generating rules analogous to those used in 1dCA (Wolfram1983, ), with memory playing the role of spatial interaction. Formally, a generating rule with memory $\mu$ is a function that maps each of the $2^{\mu}$ binary sequences of length $\mu$ to a single binary output:
$R(a_{1},a_{2},\ldots,a_{\mu})=a_{j}$ (3)
where $a_{i}\in\{0,1\}$ for all $i$ (see Fig. 1). Starting from an initial configuration, the rule is applied recursively to generate the digits of the sequence as follows:
$a_{j+1}=R(a_{j-\mu+1},a_{j-\mu+2},\ldots,a_{j})$ (4)
which eventually converges to an asymptotic pattern with period $T$.
For example, when $\mu=3$, there are $2562^{2^{3}}=$ 256 possible rules—corresponding to Wolfram’s elementary 1dCA rules. Figure 1A shows the XOR rule for $\mu=3$, represented by the binary rule string 10010110, which corresponds to rule number 150 in Wolfram’s classification. Applying this rule to the initial configuration 0 1 00100\,1\,00 1 0 yields:
[…]
which has period $T_{5}=2$.
In general, for a given $\mu$, the total number of possible generating rules is $2^{2^{\mu}}$, and the maximum number of binary sequences generated from these rules is:
$C(\mu)=2^{2^{\mu}+\mu}$
This number grows super-exponentially with $\mu$, making exhaustive analysis computationally infeasible for large $\mu$. For instance, when $\mu=10$, there are approximately $10^{308}$ rules, each of which must be applied to 1024 different initial configurations.
This raises the fundamental question: given a memory value $\mu$ and an initial configuration, is it possible to predict the asymptotic pattern and period $T$ of the resulting sequence? While this is tractable for small $\mu$ via exhaustive enumeration, alternative methods are necessary for large $\mu$ due to the sheer size of the rule space.
Figure 1: Two generating rules with memory $\mu=3$. A) Rule 150, whose binary representation is $10010110$. This corresponds to the following assignments for each 3-tuple (ordered from $111$ to $000$): $111\rightarrow 1$, $110\rightarrow 0$, $101\rightarrow 0$, $100\rightarrow 1$, $011\rightarrow 0$, $010\rightarrow 1$, $001\rightarrow 1$, $000\rightarrow 0$. The application of this rule to four different initial triplets is shown on the left. As observed, the resulting sequences are periodic with periods 1, 4, 1, and 2 (from top to bottom). The dynamics of this rule can be visualized as a directed graph, shown at the bottom. Each node represents a possible 3-tuple; a directed edge from node $x$ to node $y$ exists if applying the rule to $x$ yields the last digit of $y$. This graph is not connected and does not contain a Hamiltonian path. B) Rule 45, with binary representation $00101101$, corresponds to the truth table shown on the left of panel B. In contrast to Rule 150, applying Rule 45 to any of the $82^{3}=$ 8 initial triplets yields the same cyclic sequence of maximum period $T_{\text{max}}=8$, namely, the de Bruijn sequence $s_{6}=00010111$. This rule is thus a de Bruijn rule. The associated graph, built as described above, is a de Bruijn graph because it contains a Hamiltonian path through all $2^{3}$ nodes.
It is worth noting that the 0-dimensional Cellular Automata (0dCA) with memory, as defined above, are equivalent to sequences generated by a Linear Feedback Shift Register (LFSR). This type of sequence has been extensively studied, and a vast body of literature exists on the subject (Golomb, ). In particular, significant attention has been devoted to the characterization and computation of shift register sequences with maximum period—commonly known as de Bruijn sequences (deBruijn, ; Fredricksen, ; Ralston, ). As we will show in the next section, there exists a one-to-one correspondence between de Bruijn sequences and their generating rules, which we will refer to as de Bruijn rules.
3 Rules that generate maximum period sequences: de Bruijn rules
For a given memory length $\mu$, the maximum period that a generated binary sequence can attain is $T_{\text{max}}=2^{\mu}$. Naturally, the minimum period is 1, corresponding to sequences that converge to fixed points (e.g., either 0 or 1). For small values of $\mu$, it is feasible to generate all possible sequences and compute their corresponding periods. Figure 2 and Table 1 show, respectively, the bar chart of the observed periods and the associated frequencies for the case $\mu=4$.
Table 1: Number of rules that generates sequences with each of the possible periods that can appear, from 1 to 16, for $\mu=4$. The bar chart exhibits three peaks at $T=1$, $T=3$ and $T=5$. For $T>5$ the distribution decreases monotonously (see Fig. 2)
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
16784 7220 9547 8060 8665 5668 3494 2670 1592 977 421 192 90 77 63 16
Figure 2: Bar chart of periods for a memory $\mu=4$. The frequencies correspond to the values presented in Table 1. Similar bar charts appear for other $\mu$-values. Note that the maximun period has the lower occurrence; as a matter of fact, there are $2^{2^{3}-4}$ de Bruijn sequences.
Sequences with maximum period, known as de Bruijn sequences, are of particular interest. These are cyclic sequences of length $2^{\mu}$ in which every possible binary substring of length $\mu$ appears exactly once. For small values of $\mu$, it is straightforward to generate all possible de Bruijn sequences. There exists a unique de Bruijn sequence for $\mu=2$, namely 0011. For $\mu=3$, there are two such sequences: 00010111 and 00011101, that are generated by rules 45 and 75, respectively (rule 45 is graphically described in Figure 1B).
Since these sequences are cyclic, their representations are not unique; by convention, we select the lexicographically least sequence (i.e., the smallest in binary order and decimal number) as the representative of the equivalence class. Specially significative is the lexicographically least sequence for each memory $\mu$, that is referred to as the granddaddy in Knuth .
The number of binary de Bruijn sequences for a given $\mu$ is given by $C(\mu)=2^{2^{\mu-1}-\mu}$ (deBruijn, ; A016031, ). This number grows also superexponentially with $\mu$ and reaches extremely large values even for moderate $\mu$. For example, for $\mu=10$, $C(10)\approx 10^{151}$. Generating all de Bruijn sequences is a challenging combinatorial problem that remains unsolved in full generality, though various efficient construction algorithms have been proposed (see, for instance, deBruijnproject ).
De Bruijn sequences can be generated by applying certain specific rules, independently of the initial configuration of length $\mu$. In other words, there exists a bijection between de Bruijn sequences and their corresponding rules, which we refer to as de Bruijn rules . Table 2 lists the de Bruijn rules and their de Bruijn sequences for $\mu=1,2,3$, and $4$, expressed both as binary maps and in their corresponding decimal representations.
Table 2: Correspondence betweeen de Bruijn rules and de Bruijn sequences for $\mu=1,2,3$ and $4$. It is also shown the Evil Odd Number that divides each de Bruijn rule, leaving the remainder $\phi(\mu)$.
$\mu$ Evil odd number $\phi(\mu)$ Rule in decimal Rule in binary de Bruijn Sequence
1 - - 1 01 01
2 3 1 3 0011 0011
3 3 15 45 00101101 00010111
3 5 15 75 01001011 00011101
4 3 255 765 0000001011111101 0000101101001111
4 9 255 2295 0000100011110111 0000110100101111
4 15 255 3825 0000111011110001 0000100110101111
4 17 255 4335 0001000011101111 0000111100101101
4 27 255 6885 0001101011100101 0000101111001101
4 29 255 7395 0001110011100011 0000110101111001
4 43 255 10965 0010101011010101 0000101001101111
4 57 255 14535 0011100011000111 0000110111100101
4 65 255 16575 0100000010111111 0000111101001011
4 71 255 18105 0100011010111001 0000100111101011
4 75 255 19125 0100101010110101 0000101111010011
4 83 255 21165 0101001010101101 0000101100111101
4 85 255 21675 0101010010101011 0000111101011001
4 89 255 22695 0101100010100111 0000110010111101
4 99 255 25245 0110001010011101 0000101001111011
4 113 255 28815 0111000010001111 0000111101100101
For $\mu=4$, there are $C(4)=2^{2^{3}-4}=16$ de Bruijn rules and their associated de Bruijn sequences. As it can be seen, the granddaddy sequence is: $0000100110101111$. The corresponding granddaddy rule is $3825$, with binary representation $0000111011110001$. For $\mu=5$, the number of de Bruijn sequences is $C(5)=2^{2^{4}-5}=2048$. The granddaddy rule in this case is $00001100111111101111001100000001$ (whose decimal representation is 218034945), which generates the granddaddy sequence: $00000100011001011011101010011111$.
4 Characterization of de Bruijn rules
As stated in the previous section, certain generating rules produce sequences with maximum period—we refer to them as de Bruijn sequences. Therefore, a complete characterization of such rules emerges as a fundamental goal. The following properties help identify and constrain the set of de Bruijn rules, and can be used to systematically search for them:
  • Boundary Conditions: The binary representations of de Bruijn rules must start with 00 and end with $1$. This condition arises from the convention in rule ordering, which arranges input strings from the highest binary value ($11\ldots 1$ 11) to the lowest ($00\ldots 0$ 00), and from the necessity to avoid fixed points (e.g., if $R(11\ldots 11)=1$, the sequence remains constant). This constraint immediately reduces the number of candidate rules by a factor of 4. For example, for $\mu=4$, the number of potentially valid rules is reduced from $655362^{2^{4}}=$ 65536 to $16384$.
  • Symmetry and Parity: The binary representations of de Bruijn rules are symmetric with respect to their midpoint, such that each half is the complement of the other. This symmetry ensures parity: the number of 00s equals the number of $1$s, although the balance may be broken within each half. Under this constraint, valid de Bruijn rules correspond to multiples of certain numbers derived from sequences of Evil Odd numbers (A129771, ), multiplied by factors denoted as $\phi(\mu)$.
    For instance, if we denote the Evil Odd Numbers, i.e., an odd number whose binary representation contains an even number of $1$, for each $\mu$ as $R(\mu)_{k}$, being $k=1,2,\ldots,C(\mu)$, we get:
    $R(2)_{1}=1\times 3=3,\quad R(3)_{1}=3\times 15=45,\quad R(3)_{2}=5\times 15=75$
    The factors follow the recursive relation:
    $\phi(\mu+1)=\phi(\mu)\cdot\left(2^{2^{\mu}}+1\right)$ (5)
    For instance, $\phi(2)=1$, $\phi(3)=15$, $\phi(4)=255$, and $\phi(5)=65535$. Table 2 presents all de Bruijn rules for $\mu=4$ whose decimal representations are products of an Evil Odd number and the corresponding $\phi(\mu)$. Note, however, that not all Evil Odd numbers yield valid de Bruijn rules, and the problem of determining which ones do remains open for larger values of $\mu$. For $\mu=4$, applying these conditions reduce the feasible set of de Bruijn rules to 32.
  • Constrained Position Pairs: There exist pairs of positions in the binary representation of de Bruijn rules that cannot simultaneously take the same value (1 for even $\mu$-values and 0 for odd ones). This condition is applied only to the first half of the binary string for symmetry (according to the previous item) and depends on $\mu$ according to a recursive pattern. Specifically: For $\mu=2$, the constrained positions are $p^{1}_{2}=1$ and $p^{2}_{2}=2\,p^{1}_{2}$, and
    • If $\mu=2k+1$, then $p^{1}_{2k+1}=p^{2}_{2k}$ and $p^{2}_{2k+1}=2\,p^{1}_{2k+1}-1$.
    • If $\mu=4k$, then $p^{1}_{4k}=p^{2}_{4k-1}$ and $p^{2}_{4k}=2\,p^{1}_{4k}$,
    for $k=1,2,\ldots$. This structural condition eliminates $1/4$ of the remaining candidates. For example, for $\mu=4$ these positions are $p^{1}_{4}=3$ and $p^{2}_{4}=6$. For $\mu=5$, $p^{1}_{5}=6$ and $p^{2}_{5}=11$.
    Remarkably, for $\mu=4$, the final count of feasible de Bruijn rules after applying all constraints is 24. In general, for all values of $\mu$, the feasible set exceeds the actual number of de Bruijn rules (see Table 3).
  • Symmetric Rule Invariance: If a rule of the form:
[…]
    is a de Bruijn rule, then its mirrored version:
[…]
    is also a de Bruijn rule.
    This property reflects the inherent symmetry and reversibility in de Bruijn rule structure. It ensures that for each valid de Bruijn rule constructed in this way, a corresponding reverse-complement rule also exists within the de Bruijn set.
The application of the properties stated in the previous paragraphs to the entire set of generating rules significantly reduces the number of feasible rules that can yield de Bruijn sequences. Table 3 presents the total number of rules, the number of de Bruijn sequences, the number of feasible rules after applying the constraints, and the corresponding ratios.
As can be observed, the number of feasible rules that need to be checked to find de Bruijn rules is drastically smaller than the total number of rules for each $\mu$ value. For instance, when $\mu=6$, the reduction factor is approximately $10^{-11}$. In the same case, the ratio between the number of rules that actually yield de Bruijn sequences and the number of feasible rules is around $0.17$. For such low values of $\mu$, a brute-force approach may still be used to generate all de Bruijn sequences. However, a more efficient and structured methodology will be presented in the next section.
Table 3: Table resulting from the application of the properties described in Section 4. The second column, $C(\mu)$, indicates the total number of rules for each value of $\mu$. The third column shows the number of feasible rules remaining after applying the constraints detailed in Section 4. The fourth column lists the number of de Bruijn rules. The remaining columns present ratios between these subsets. Particularly noteworthy is the fifth column, which displays the ratio between the feasible subset and the total number of rules. As $\mu$ increases, the reduction in the search space for de Bruijn rules becomes dramatic — for example, for $\mu=9$, the feasible subset constitutes only about $10^{-79}$ of the full rule set.
$\mu$ $C(\mu)$#\## Feasible ##\## de Bruijn Feasible/Total de Bruijn/Total de Bruijn/Feasible
2 16 1 0.0625
3 256 2 2 0.0078125 0.0078125 1
4 65536 24 16 0.000366211 0.000244141 0.66666667
5 4294967296 6144 2048 1.4305E-06 4.7683E-07 0.33333333
6 1.84467E+19 402653184 67108864 2.1827E-11 3.6379E-12 0.16666667
7 3.40282E+38 1.7293E+18 1.4411E+17 5.082E-21 4.2351E-22 0.08333333
8 1.15792E+77 3.1901E+37 1.3292E+36 2.7550E-40 1.1479E-41 0.04166667
9 1.3408E+154 1.0855E+76 2.2615E+74 8.0964E-79 1.6867E-80 0.02083333
5 Neural Networks to classify de Bruijn rules
An alternative approach to identifying de Bruijn rules is to apply machine learning methods. In particular, neural networks are especially well-suited for classification tasks Bishop . In this section, we present a neural network model for classifying the feasible rules into two categories: de Bruijn rules (coded by 1) and the rest (0). Although classification based on the sequence period is also possible, it would require accounting for dependence on initial conditions. The classification is performed for $\mu=5$ and $\mu=6$ (see Table 4).
We implemented a binary classification model using a feedforward neural network in R Rcore . The analysis followed these main steps:
  1. 1.
    Data loading and preprocessing: The input data consisted of a character string representing a binary sequence and a binary integer label.
  2. 2.
[…]
  3. 3.
    Dataset splitting: The data was randomly split into training (80%) and testing (20%) subsets to evaluate model performance on unseen data.
  4. 4.
    Model specification and training: A feedforward neural network was constructed using the keras package with a TensorFlow backend Allaire ; Rcore .
Feedforward neural networks are highly suitable for structured, tabular data, such as the bit vectors extracted from rule representations. They can capture complex, non-linear relationships between input features without requiring explicit feature engineering. In our case, the binary inputs represent discrete features, and their interactions are not trivially captured by simpler linear models. The multi-layer architecture enables hierarchical feature learning, significantly improving classification accuracy.
For the case $\mu=5$, the neural network architecture includes:
  • An input layer with 14 features (bits),
  • A hidden dense layer with 32 units and ReLU activation,
  • A second hidden dense layer with 16 units and ReLU activation,
  • An output layer with 1 unit and sigmoid activation for binary classification.
The model was compiled using the Adam optimizer (learning rate $=0.001$), binary cross-entropy loss function, and accuracy as a performance metric. The dataset includes the complete set of 6144 feasible rules, one-third of which are de Bruijn rules (see Table 3). Training was performed over 100 epochs with a batch size of 4. Model evaluation was conducted on the test set by calculating accuracy, sensitivity, and specificity (see Table 4). Class labels were assigned using a threshold of 0.5 on the predicted probabilities. The resulting classifier achieved outstanding performance, with an accuracy exceeding 99% for $\mu=5$.
The more challenging case of $\mu=6$ involves a rule space of size on the order of $10^{19}$, with approximately 67 million de Bruijn rules (see Table 3). As a matter of fact, we can only use a sample of the total rule space that is randomly analyzed and classified into the two classes.
Optimal results were achieved using a slightly deeper neural network with the following architecture:
  • Three hidden dense layers with 64, 64, and 8 units respectively,
  • ReLU activations in all hidden layers,
  • A sigmoid output unit for binary classification.
Training used a batch size of 64 and a learning rate of $0.001$. The dataset consists of a sample of $2\times 10^{6}$ feasible rules, half of which are de Bruijn rules. As before, 80% of the dataset was used for training, and the remainder for testing. A threshold of 0.5 was again applied to the output probabilities to assign class labels. As shown in Table 4, this model also demonstrated excellent classification metrics.
Table 4: Evaluation metrics for the neural network models applied to classify the de Bruijn rules for memories with $\mu=5$ and $\mu=6$. Total refers to the sum: $TP+FP+TN+FN$, and corresponds to one fith of the whole data set. Positive class considered: 1 $\to$ de Bruijn.
Metric (definition) $\mu=5$ $\mu=6$
True Positives (TP) 397 198563
False Positives (FP) 3 19839
True Negatives (TN) 820 180668
False Negatives (FN) 9 930
Accuracy, (TP + TN) / Total 0.9902 0.9481
Sensitivity (Recall), TP / (TP + FN) 0.9780 0.9953
Specificity, TN / (TN + FP) 0.9964 0.9011
Precision (PPV), TP / (TP + FP) 0.9925 0.9092
Negative Predictive Value (NPV), TN / (TN + FN) 0.9891 0.9949
Balanced Accuracy, (Sens. + Spec.) / 2 0.9872 0.9482
Detection Rate, TP / Total 0.3230 0.4976
Detection Prevalence, (TP + FP) / Total 0.3255 0.5473
True Prevalence, (TP + FN) / Total 0.3304 0.4999
Once the neural network model is available, any given rule can be evaluated to determine, with high probability, whether it is a de Bruijn rule. If the prediction is positive, the rule is then applied to generate the corresponding binary sequence, which must ultimately be verified to confirm that it satisfies the de Bruijn sequence properties.
6 Discussion
The successive application of updating rules to an initial configuration of $\{0,1\}$ generates a binary sequence that becomes asymptotically periodic. The size of the initial configuration depends on the memory of the rule, denoted by $\mu$, which is the number of digits required to update the next one. For large values of $\mu$, the number of possible generating rules becomes so large that an exhaustive analysis of the resulting patterns is infeasible. Particularly relevant is a specific and highly constrained subset of rules, that we named as de Bruijn rules, that generate sequences of maximum period, known as de Bruijn sequences. Although these sequences represent only a tiny fraction of the entire set of possible sequences, their sheer number for large $\mu$ remains enormous, and the complete and effective generation of all such sequences is still an open problem.
In this paper, we have presented a novel approach to compute de Bruijn sequences that combines two complementary methodologies. First, by exploiting structural properties of de Bruijn rules—i.e., those rules that generate maximum period sequences—a drastic reduction of the full set of candidate rules can be achieved. Table 3 summarizes the reduction ratios for several values of $\mu$. Second, we applied a machine learning approach to this feasible subset in order to accurately identify the de Bruijn rules. As shown in Section 5, the use of a classical neural network model for $\mu=5$, and $6$ allows for nearly complete classification of de Bruijn rules in the smaller cases, and achieves over 99% and 94% accuracy for $\mu=5$ and $\mu=6$, respectively. Once the de Bruijn rules are identified, the corresponding de Bruijn sequences are straightforward to construct. This would allow to find granddaddy sequence that represents the lexicographically smallest member for each $\mu$-value.
To the best of our knowledge, this is the first attempt to generate de Bruijn sequences using a hybrid approach that integrates Machine Learning techniques within the framework of Cellular Automata for generating maximum period sequences. It is worth mentioning that existing methods based on Feedback Shift Register generation have not definitively solved the problem. Therefore, we believe that the results presented in this paper represent a significant step forward in the study and generation of this important class of sequences.

### v2

@@
 We investigate binary sequences generated by non-Markovian rules with memory length $\mu$, similar to those adopted in elementary cellular automata.
-This generation procedure is equivalente to a shift register and certain rules produce sequences with maximal periods, known as de Bruijn sequences.
-We introduce a novel methodology for generating de Bruijn sequences that combines: (i) a set of derived properties that significantly reduce the space of feasible generating rules, and (ii) a neural network-based classifier that identifies which rules produce de Bruijn sequences.
-Experiments for large values of $\mu$ demonstrate the approach’s effectiveness and computational efficiency.
+This generation procedure is equivalent to a shift register, and certain rules produce sequences with maximal periods, known as de Bruijn sequences.
+We introduce a novel methodology for generating de Bruijn sequences that combines (i) a set of derived properties that significantly reduce the space of feasible generating rules and (ii) a neural-network-based classifier that identifies which rules produce de Bruijn sequences.
+The experiments for some values of $\mu$ demonstrate the approach’s effectiveness and computational efficiency.
 1 Introduction
@@
 The CA evolves by applying generating rules at each time step.
-These rules can be static or dynamic (i.e., they may change depending on the current state of the entire CA), and either local or global, depending on whether they involve only neighboring cells or the entire population.
+These rules can be static or dynamic (i.e., they may change depending on the current state of the entire CA) and either local or global, depending on whether they involve only neighboring cells or the entire population.
 From an initial configuration, the state of the CA evolves over time according to the specified rules.
@@
 Generally, the complexity of the system increases with both $\mu$ and $\alpha$.
-For simplicity, we focus on binary systems, i.e., $\alpha=2$, using the symbols $\{0,1\}$.
+For simplicity, we focus on binary systems, i.e., $\alpha=2$, using the symbols $\{0,1\}$, and, due to computational limitations, we restrict our study to a subset of cases with relatively low values of $\mu$ (less than 8).
 2 0-Dimensional CA
-Symbolic sequences naturally arise in discrete dynamical systems, e.g.
-CA (Jin, ).
+Symbolic sequences naturally arise in discrete dynamical systems, e.g., CA (Jin, ).
 From an initial condition, the next state is computed according to a generation rule (or function) that depends on previous states.
-For instance, a classical non-Markovian binary time series can be defined by initial values $a_{1}$ and $a_{2}$, and a recursive rule:
+For instance, a classical non-Markovian binary time series can be defined by initial values $a_{1}$ and $a_{2}$ and a recursive rule:
 $a_{k+2}=(a_{k+1}+a_{k})\mod 2,\quad\text{for }k=1,2,\ldots$ (1)
-With initial states $a_{1}=0$ and $a_{2}=1$, the sequence continues as $a_{3}=1$, $a_{4}=0$, $a_{5}=1$, $a_{6}=1$, etc., producing the binary sequence:
+With initial states $a_{1}=0$ and $a_{2}=1$, the sequence continues as $a_{3}=1$, $a_{4}=0$, $a_{5}=1$, $a_{6}=1$, etc., producing the following binary sequence:
 […]
 This sequence is clearly periodic with period $T_{1}=3$. In this case, the next digit depends on the two previous digits, implying a memory $\mu=2$.
-Other examples with $\mu=2$ can be constructed using logical operators. For instance, applying the AND function with the rules:
-1 11,1 00,0 10,0 00formulae-sequence111formulae-sequence100formulae-sequence0100001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 0$\quad$ 0\,1$\rightarrow$ 0$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 0 , 0 1 → 0 , 0 0 → 0
-to the initial pair 0 1010\,10 1 produces the sequence:
+Other examples with $\mu=2$ can be constructed using logical operators. For instance, applying the AND-function with the rules
 […]
-Similarly, applying the OR function:
-1 11,1 01,0 11,0 00formulae-sequence111formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 1$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 1 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
-to the same initial configuration 0 1010\,10 1 results in:
-[…]
+to the initial pair 0 10\,1 produces the sequence
+$s_{2}=0\,1\,0\,0\,0\,0\,0\,\ldots$
+Similarly, applying the OR-function
+1 11,1 $\rightarrow 1,\quad 1\,0\rightarrow 1,\quad 0\,1\rightarrow 1,\quad 0\,0\rightarrow 0$
+to the same initial configuration 0 10\,1 results in
+$s_{3}=0\,1\,1\,1\,1\,1\,1\,\ldots$
 Both sequences converge to fixed values (0 or 1), i.e., their periods are $T_{2}=T_{3}=1$, and these fixed points are independent of the initial configuration.
-In contrast, applying the XOR function:
-1 10,1 01,0 11,0 00formulae-sequence110formulae-sequence101formulae-sequence0110001\,1$\rightarrow$ 0$\quad$ 1\,0$\rightarrow$ 1$\quad$ 0\,1$\rightarrow$ 1$\quad$ 0\,0$\rightarrow$ 01 1 → 0 , 1 0 → 1 , 0 1 → 1 , 0 0 → 0
-to the initial state $a_{1}=0,a_{2}=1$ yields:
+In contrast, applying the XOR-function
+1 10,1 01,0 11,0 $\rightarrow 0,\quad 1\,0\rightarrow 1$$\quad$ $\rightarrow 1,\quad 0\,$$\rightarrow 0$
+to the initial state $a_{1}=0,a_{2}=1$ yields
 […]
 with period $T_{4}=3$.
-However, using the same rule with the initial configuration: $a_{1}=0,a_{2}=0$ results in a constant sequence $a_{k}=0$ for all $k>2$.
+However, using the same rule with the initial configuration $a_{1}=0,$ $a_{2}=0$ results in a constant sequence $a_{k}=0$ for all $k>2$.
 These examples illustrate that the resulting behavior depends not only on the rule but also on the initial conditions.
@@
 That is, for any rule and initial condition, the sequence will eventually settle into a repeating pattern (also referred to as a motif or loop).
-The maximum possible period for a sequence with memory $\mu$ is:
+The maximum possible period for a sequence with memory $\mu$ is
 $T_{\text{max}}=2^{\mu}$ (2)
-As $\mu$ increases, the sequences may resemble aperiodic ones, making them useful for computing pseudorandom numbers to be applied in Monte Carlo simulations, and cryptographic systems (Etzion, ).
+As $\mu$ increases, the sequences may resemble aperiodic ones, making them useful for computing pseudorandom numbers to be applied in Monte Carlo simulations and cryptographic systems (Etzion, ).
 Truly aperiodic behavior can only arise when $\mu\rightarrow\infty$, or when the alphabet is infinite (e.g., using real-valued states).
@@
 $R(a_{1},a_{2},\ldots,a_{\mu})=a_{j}$ (3)
-where $a_{i}\in\{0,1\}$ for all $i$ (see Fig. 1). Starting from an initial configuration, the rule is applied recursively to generate the digits of the sequence as follows:
+where $a_{i}\in\{0,1\}$ for all $i=1,2,\ldots,\mu$ (see Figure 1). Starting from an initial configuration, the rule is applied recursively to generate the digits of the sequence as follows:
 $a_{j+1}=R(a_{j-\mu+1},a_{j-\mu+2},\ldots,a_{j})$ (4)
 which eventually converges to an asymptotic pattern with period $T$.
-For example, when $\mu=3$, there are $2562^{2^{3}}=$ 256 possible rules—corresponding to Wolfram’s elementary 1dCA rules.
-Figure 1A shows the XOR rule for $\mu=3$, represented by the binary rule string 10010110, which corresponds to rule number 150 in Wolfram’s classification.
-Applying this rule to the initial configuration 0 1 00100\,1\,00 1 0 yields:
-[…]
-which has period $T_{5}=2$.
-In general, for a given $\mu$, the total number of possible generating rules is $2^{2^{\mu}}$, and the maximum number of binary sequences generated from these rules is:
-$C(\mu)=2^{2^{\mu}+\mu}$
-This number grows super-exponentially with $\mu$, making exhaustive analysis computationally infeasible for large $\mu$.
-For instance, when $\mu=10$, there are approximately $10^{308}$ rules, each of which must be applied to 1024 different initial configurations.
-This raises the fundamental question: given a memory value $\mu$ and an initial configuration, is it possible to predict the asymptotic pattern and period $T$ of the resulting sequence?
-While this is tractable for small $\mu$ via exhaustive enumeration, alternative methods are necessary for large $\mu$ due to the sheer size of the rule space.
 Figure 1: Two generating rules with memory $\mu=3$.
-A) Rule 150, whose binary representation is $10010110$.
-This corresponds to the following assignments for each 3-tuple (ordered from $111$ to $000$): $111\rightarrow 1$, $110\rightarrow 0$, $101\rightarrow 0$, $100\rightarrow 1$, $011\rightarrow 0$, $010\rightarrow 1$, $001\rightarrow 1$, $000\rightarrow 0$.
-The application of this rule to four different initial triplets is shown on the left.
+(A) Rule 150, whose binary representation is $10010110$.
+This corresponds to the following assignments for each 3-tuple (ordered from $111$ to $000$): $11\rightarrow 1, 11$$\rightarrow 0, 1010$$1\rightarrow 0$, $0\rightarrow 1, 0110$$1\rightarrow 0$, $\rightarrow 1, 001100$$\rightarrow 1$, $000\rightarrow 0$.
+The application of this rule to four different initial triplets is shown on the right of panel (A).
 As observed, the resulting sequences are periodic with periods 1, 4, 1, and 2 (from top to bottom).
@@
 This graph is not connected and does not contain a Hamiltonian path.
-B) Rule 45, with binary representation $00101101$, corresponds to the truth table shown on the left of panel B.
-In contrast to Rule 150, applying Rule 45 to any of the $82^{3}=$ 8 initial triplets yields the same cyclic sequence of maximum period $T_{\text{max}}=8$, namely, the de Bruijn sequence $s_{6}=00010111$.
+(B) Rule 45, with binary representation $00101101$, corresponds to the truth table shown on the left of panel B.
+In contrast to Rule 150, applying Rule 45 to any of the $2^{3}=8$ initial triplets yields the same cyclic sequence of maximum period $T_{\text{max}}=8$, namely, the de Bruijn sequence $s_{6}=00010111$.
 This rule is thus a de Bruijn rule.
 The associated graph, built as described above, is a de Bruijn graph because it contains a Hamiltonian path through all $2^{3}$ nodes.
-It is worth noting that the 0-dimensional Cellular Automata (0dCA) with memory, as defined above, are equivalent to sequences generated by a Linear Feedback Shift Register (LFSR).
-This type of sequence has been extensively studied, and a vast body of literature exists on the subject (Golomb, ).
-In particular, significant attention has been devoted to the characterization and computation of shift register sequences with maximum period—commonly known as de Bruijn sequences (deBruijn, ; Fredricksen, ; Ralston, ).
+For example, when $\mu=3$, there are $2^{2^{3}}=256$ possible rules—corresponding to Wolfram’s elementary 1dCA rules.
+Figure 1A shows the XOR rule for $\mu=3$, represented by the binary rule string $10010110$, which corresponds to rule number 150 in Wolfram’s classification.
+Applying this rule to the initial configuration 0 1 00\,1\,0 yields
+[…]
+which has period $T_{5}=2$.
+In general, for a given $\mu=1,2,3,\ldots$, the total number of possible generating rules is $C(\mu)=2^{2^{\mu}}$ (this number corresponds with the Fermat number minus 1 A000215 ), and the maximum number of binary sequences generated from these rules is $2^{2^{\mu}+\mu}$.
+This number grows super-exponentially with $\mu$, making exhaustive analysis computationally infeasible for large $\mu$.
+For instance, when $\mu=10$, there are approximately $10^{308}$ rules, each of which must be applied to 1024 different initial configurations.
+This raises the following fundamental question: given a memory value $\mu$ and an initial configuration, is it possible to predict the asymptotic pattern and period $T$ of the resulting sequence?
+While this is tractable for small $\mu$ via exhaustive enumeration, alternative methods are necessary for large $\mu$ due to the sheer size of the rule space.
+It is worth noting that the 0-dimensional cellular automata (0dCA) with memory, as defined above, are equivalent to sequences generated by a Non-Linear Feedback Shift Register (NLFSR).
+This type of sequence has been extensively studied, and a vast body of literature exists on the subject Golomb ; Fredricksen ; Gao ; Zhao .
+In particular, significant attention has been devoted to the characterization and computation of shift register sequences with maximum period—commonly known as de Bruijn sequences deBruijn ; Ralston ; Huang ; Miros .
 As we will show in the next section, there exists a one-to-one correspondence between de Bruijn sequences and their generating rules, which we will refer to as de Bruijn rules.
@@
 For small values of $\mu$, it is feasible to generate all possible sequences and compute their corresponding periods.
-Figure 2 and Table 1 show, respectively, the bar chart of the observed periods and the associated frequencies for the case $\mu=4$.
-Table 1: Number of rules that generates sequences with each of the possible periods that can appear, from 1 to 16, for $\mu=4$.
-The bar chart exhibits three peaks at $T=1$, $T=3$ and $T=5$.
-For $T>5$ the distribution decreases monotonously (see Fig.
-2)
-1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
-16784 7220 9547 8060 8665 5668 3494 2670 1592 977 421 192 90 77 63 16
-Figure 2: Bar chart of periods for a memory $\mu=4$.
-The frequencies correspond to the values presented in Table 1.
-Similar bar charts appear for other $\mu$-values.
-Note that the maximun period has the lower occurrence; as a matter of fact, there are $2^{2^{3}-4}$ de Bruijn sequences.
-Sequences with maximum period, known as de Bruijn sequences, are of particular interest.
+For $\mu=4$, Table 1 shows the number of rules that generate sequences with periods in between 1 and $2^{4}$ for each of the $2^{4}$ initial sequence of length $4$.
+Note that the distributions of periods are symmetric to middle rows and that the maximum period occurs for only 16 rules, independently of the initial sequence.
+The proportion of rules as a function of the maximum period of the sequences that each of the $2^{2^{4}}$ generates over all the initial sequences is depicted in Figure 2 (also for the case $\mu=4$).
+It is worthy to remark the nonmonotonous shape of this distribution that exhibits three relative maxima at periods $T=1$, $T=3$, and $T=5$.
+For $T>5$, it seems to decrease exponentially.
+Sequences with a maximum period, known as de Bruijn sequences, are of particular interest.
 These are cyclic sequences of length $2^{\mu}$ in which every possible binary substring of length $\mu$ appears exactly once.
 For small values of $\mu$, it is straightforward to generate all possible de Bruijn sequences.
-There exists a unique de Bruijn sequence for $\mu=2$, namely 0011.
-For $\mu=3$, there are two such sequences: 00010111 and 00011101, that are generated by rules 45 and 75, respectively (rule 45 is graphically described in Figure 1B).
+There exists a unique de Bruijn sequence for $\mu=2$, namely, $0011$.
+For $\mu=3$, there are two such sequences: $00010111$ and $00011101$ that are generated by rules 45 and 75, respectively (rule 45 is graphically described in Figure 1B).
 Since these sequences are cyclic, their representations are not unique; by convention, we select the lexicographically least sequence (i.e., the smallest in binary order and decimal number) as the representative of the equivalence class.
-Specially significative is the lexicographically least sequence for each memory $\mu$, that is referred to as the granddaddy in Knuth .
+Specially significative is the lexicographically least sequence for each memory $\mu$ that is referred to as the granddaddy in Knuth .
+Table 1: Number of rules that generates sequences with each of the possible periods (columns) that can appear, from 1 to 16, for $\mu=4$, applied to the $2^{4}$ initial sequences (rows).
+Notice that the total number of rules for this $\mu$-value is $2^{2^{4}}$ = 65,536, which equal the sum of each row.
+As it can be observed, the distributions of periods are symmetric to middle rows.
+Note that the maximum period occurs at the last column ($T=16$) for 16 rules and for every initial condition.
+Initial Sequence 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
+$00$ 36,096 3304 4944 4734 5836 3758 2472 1846 1120 758 332 144 80 48 48 16
+0001 12,912 6608 9888 9468 9624 6492 3920 2924 1728 1068 440 208 96 80 64 16
+0010 11,800 8976 11,780 8457 8976 5697 3552 2803 1634 995 412 200 94 80 64 16
+0011 15,368 5320 9490 9252 9296 6397 3702 2981 1756 1067 444 207 96 80 64 16
+0100 14,568 5524 11,588 8860 9190 5851 3570 2834 1672 1011 414 200 94 80 64 16
+$01$ 10,520 16,384 8458 6494 8542 5755 3620 2572 1438 919 406 192 80 76 64 16
+0110 12,032 6428 12,398 9128 9438 6043 3486 2975 1714 1019 422 199 94 80 64 16
+0111 21,376 4956 7416 8125 8242 5637 3580 2641 1648 1025 426 208 96 80 64 16
+1000 21,376 4956 7416 8125 8242 5637 3580 2641 1648 1025 426 208 96 80 64 16
+1001 12,032 6428 12,398 9128 9438 6043 3486 2975 1714 1019 422 199 94 80 64 16
+$10$ 10,520 16,384 8458 6494 8542 5755 3620 2572 1438 919 406 192 80 76 64 16
+1011 14,568 5524 11,588 8860 9190 5851 3570 2834 1672 1011 414 200 94 80 64 16
+1100 15,368 5320 9490 9252 9296 6397 3702 2981 1756 1067 444 207 96 80 64 16
+1101 11,800 8976 11,780 8457 8976 5697 3552 2803 1634 995 412 200 94 80 64 16
+1110 12,912 6608 9888 9468 9624 6492 3920 2924 1728 1068 440 208 96 80 64 16
+$11$ 36,096 3304 4944 4734 5836 3758 2472 1846 1120 758 332 144 80 48 48 16
+Figure 2: Bar chart of the proportion of rules as a function of the maximum period of the sequences that each of them generates over the initial sequences for a memory $\mu=4$.
+Note that $T=16$ has the lower occurrence; as a matter of fact, there are $2^{2^{3}-4}$ de Bruijn rules, which result in a relative frequency of $16/65,536\approx 0.0002$.
 The number of binary de Bruijn sequences for a given $\mu$ is given by $C(\mu)=2^{2^{\mu-1}-\mu}$ (deBruijn, ; A016031, ).
@@
 For example, for $\mu=10$, $C(10)\approx 10^{151}$.
-Generating all de Bruijn sequences is a challenging combinatorial problem that remains unsolved in full generality, though various efficient construction algorithms have been proposed (see, for instance, deBruijnproject ).
+Generating all de Bruijn sequences is a challenging combinatorial problem that remains unsolved in full generality, though various efficient construction algorithms have been proposed (see, for instance, Hall ; deBruijnproject ).
 De Bruijn sequences can be generated by applying certain specific rules, independently of the initial configuration of length $\mu$.
@@
 Table 2 lists the de Bruijn rules and their de Bruijn sequences for $\mu=1,2,3$, and $4$, expressed both as binary maps and in their corresponding decimal representations.
+For $\mu=4$, there are $C(4)=2^{2^{3}-4}=16$ de Bruijn rules and their associated de Bruijn sequences.
+As it can be seen, the granddaddy sequence is $0000100110101111$.
+The corresponding granddaddy rule is $3825$, with binary representation $0000111011110001$.
+For $\mu=5$, the number of de Bruijn sequences is $C(5)=2^{2^{4}-5}=2048$.
+The granddaddy rule in this case is $00001100111111101111001100000001$ (whose decimal representation is 218,034,945), which generates the granddaddy sequence $00000100011001011011101010011111$.
 Table 2: Correspondence betweeen de Bruijn rules and de Bruijn sequences for $\mu=1,2,3$ and $4$.
-It is also shown the Evil Odd Number that divides each de Bruijn rule, leaving the remainder $\phi(\mu)$.
+The table also shows the Evil Odd Number that divides each de Bruijn rule, leaving the remainder $\phi(\mu)$.
+The Evil Odd Numbers that correspond to de Bruijn rules for $\mu=5$ are presented in Appendix B.
 $\mu$ Evil Odd Number $\phi(\mu)$ Rule in Decimal Rule in Binary de Bruijn Sequence
 1 - - 1 01 01
-2 3 1 3 0011 0011
+2 - 3 3 0011 0011
 3 3 15 45 00101101 00010111
@@
 4 29 255 7395 0001110011100011 0000110101111001
-4 43 255 10965 0010101011010101 0000101001101111
-4 57 255 14535 0011100011000111 0000110111100101
-4 65 255 16575 0100000010111111 0000111101001011
-4 71 255 18105 0100011010111001 0000100111101011
-4 75 255 19125 0100101010110101 0000101111010011
-4 83 255 21165 0101001010101101 0000101100111101
-4 85 255 21675 0101010010101011 0000111101011001
-4 89 255 22695 0101100010100111 0000110010111101
-4 99 255 25245 0110001010011101 0000101001111011
-4 113 255 28815 0111000010001111 0000111101100101
-For $\mu=4$, there are $C(4)=2^{2^{3}-4}=16$ de Bruijn rules and their associated de Bruijn sequences.
-As it can be seen, the granddaddy sequence is: $0000100110101111$.
-The corresponding granddaddy rule is $3825$, with binary representation $0000111011110001$.
-For $\mu=5$, the number of de Bruijn sequences is $C(5)=2^{2^{4}-5}=2048$.
-The granddaddy rule in this case is $00001100111111101111001100000001$ (whose decimal representation is 218034945), which generates the granddaddy sequence: $00000100011001011011101010011111$.
+4 43 255 10,965 0010101011010101 0000101001101111
+4 57 255 14,535 0011100011000111 0000110111100101
+4 65 255 16,575 0100000010111111 0000111101001011
+4 71 255 18,105 0100011010111001 0000100111101011
+4 75 255 19,125 0100101010110101 0000101111010011
+4 83 255 21,165 0101001010101101 0000101100111101
+4 85 255 21,675 0101010010101011 0000111101011001
+4 89 255 22,695 0101100010100111 0000110010111101
+4 99 255 25,245 0110001010011101 0000101001111011
+4 113 255 28,815 0111000010001111 0000111101100101
 4 Characterization of de Bruijn Rules
-As stated in the previous section, certain generating rules produce sequences with maximum period—we refer to them as de Bruijn sequences.
+As stated in the previous section, certain generating rules produce sequences with maximum period—we refer to them as de Bruijn rules.
 Therefore, a complete characterization of such rules emerges as a fundamental goal.
-The following properties help identify and constrain the set of de Bruijn rules, and can be used to systematically search for them:
-  • Boundary Conditions: The binary representations of de Bruijn rules must start with 00 and end with $1$.
-This condition arises from the convention in rule ordering, which arranges input strings from the highest binary value ($11\ldots 1$ 11) to the lowest ($00\ldots 0$ 00), and from the necessity to avoid fixed points (e.g., if $R(11\ldots 11)=1$, the sequence remains constant).
+The following properties help identify and constrain the set of de Bruijn rules and can be used to systematically search for them:
+  • Boundary Conditions: The binary representations of de Bruijn rules must start with 0 and end with $1$.
+This condition arises from the convention in rule ordering, which arranges input strings from the highest binary value ($11\ldots 11$) to the lowest ($00\ldots 00$) and from the necessity to avoid fixed points (e.g., if $R(11\ldots 11)=1$, the sequence remains constant).
 This constraint immediately reduces the number of candidate rules by a factor of 4.
-For example, for $\mu=4$, the number of potentially valid rules is reduced from $655362^{2^{4}}=$ 65536 to $16384$.
+For example, for $\mu=4$, the number of potentially valid rules is reduced from $2^{2^{4}}$ = 65,536 to 16,384.
   • Symmetry and Parity: The binary representations of de Bruijn rules are symmetric with respect to their midpoint, such that each half is the complement of the other.
-This symmetry ensures parity: the number of 00s equals the number of $1$s, although the balance may be broken within each half.
-Under this constraint, valid de Bruijn rules correspond to multiples of certain numbers derived from sequences of Evil Odd numbers (A129771, ), multiplied by factors denoted as $\phi(\mu)$.
-    For instance, if we denote the Evil Odd Numbers, i.e., an odd number whose binary representation contains an even number of $1$, for each $\mu$ as $R(\mu)_{k}$, being $k=1,2,\ldots,C(\mu)$, we get:
-    $R(2)_{1}=1\times 3=3,\quad R(3)_{1}=3\times 15=45,\quad R(3)_{2}=5\times 15=75$
-    The factors follow the recursive relation:
-    $\phi(\mu+1)=\phi(\mu)\cdot\left(2^{2^{\mu}}+1\right)$ (5)
-    For instance, $\phi(2)=1$, $\phi(3)=15$, $\phi(4)=255$, and $\phi(5)=65535$.
+This symmetry ensures parity: the number of 0s equals the number of $1$s, although the balance may be broken within each half.
+Under this constraint, valid de Bruijn rules correspond to multiples of certain Numbers derived from sequences of Evil Odd numbers (A129771, ), multiplied by factors denoted as $\phi(\mu)$ (see Appendix A).
+Let $R_{\mu}$ be a rule for a given $\mu$.
+Then,
+    $R_{\mu}=\phi(\mu)\,(A+1)$ (5)
+    In Appendix A, we show that the factor $\phi(\mu)$ is
+    $\phi(\mu)=2^{2^{\mu-1}}-1$ (6)
+    For instance, $\phi(2)=3$, $\phi(3)=15$, $\phi(4)=255$, and $\phi(5)$ = 65,535.
+The consideration of this property causes a huge reduction in the feasible set of the de Bruijn rules, from 16,384 to 64 for the case $\mu=4$.
+  • Evil Odd Number divisibility: It is also empirical evidence that factor $A+1$ in Equation (5) must be an Evil Odd Number for $R_{\mu}$ to be a de Bruijn rule.
+For each $\mu$, there is a unique $\phi$ that divides $R_{\mu}$, which results in multiple remainders that are Evil Odd Numbers.
+To consider that, it is convenient to denote this factor as $A(\mu)_{k}$, representing the Evil Odd Numbers for $k=1,2,\ldots,C(\mu)$.
+For instance,
+    $A(2,1)=1\times 3=3,\quad A(3,1)=3\times 15=45,\quad A(3,2)=5\times 15=75$
     Table 2 presents all de Bruijn rules for $\mu=4$ whose decimal representations are products of an Evil Odd Number and the corresponding $\phi(\mu)$.
@@
 This condition is applied only to the first half of the binary string for symmetry (according to the previous item) and depends on $\mu$ according to a recursive pattern.
-Specifically: For $\mu=2$, the constrained positions are $p^{1}_{2}=1$ and $p^{2}_{2}=2\,p^{1}_{2}$, and
+    Let $p_{\mu}^{j}$ be the binary value of the $j$ position in a de Bruijn rule of memory $\mu$. For $\mu=2$, the constrained positions are $p^{1}_{2}=1$ and $p^{2}_{2}=2\,p^{1}_{2}$, and
     • If $\mu=2k+1$, then $p^{1}_{2k+1}=p^{2}_{2k}$ and $p^{2}_{2k+1}=2\,p^{1}_{2k+1}-1$.
-    • If $\mu=4k$, then $p^{1}_{4k}=p^{2}_{4k-1}$ and $p^{2}_{4k}=2\,p^{1}_{4k}$,
+    • If $\mu=2\,(k+1)$, then $p^{1}_{2\,(k+1)}=p^{2}_{2\,k+1}$ and $p^{2}_{2\,(k+1)}=2\,p^{1}_{2\,(k+1)}$.
     for $k=1,2,\ldots$.
     This structural condition eliminates $1/4$ of the remaining candidates.
-For example, for $\mu=4$ these positions are $p^{1}_{4}=3$ and $p^{2}_{4}=6$.
-For $\mu=5$, $p^{1}_{5}=6$ and $p^{2}_{5}=11$.
+For example, for $\mu=3$, these positions are $p^{1}_{3}=2$ and $p^{2}_{3}=3$.
+For $\mu=4$, $p^{1}_{4}=3$ and $p^{2}_{4}=6$, and for $\mu=5$, $p^{1}_{5}=6$ and $p^{2}_{5}=11$.
     Remarkably, for $\mu=4$, the final count of feasible de Bruijn rules after applying all constraints is 24.
-In general, for all values of $\mu$, the feasible set exceeds the actual number of de Bruijn rules (see Table 3).
-  • Symmetric Rule Invariance: If a rule of the form:
+In general, for all values of $\mu>3$, the feasible set exceeds the actual number of de Bruijn rules (see Table 3).
+  • Symmetric Rule Invariance: If a rule of the form
 […]
-    is a de Bruijn rule, then its mirrored version:
-[…]
+    is a de Bruijn rule, then its mirrored version
+    $0\,a_{k}\,a_{k-1}\,\ldots\,a_{2}\,a_{1}\,0\,1\,1-a_{k}\,1-a_{k-1}\,\ldots\,1-a_{2}\,1-a_{1}\,1$
     is also a de Bruijn rule.
@@
 It ensures that for each valid de Bruijn rule constructed in this way, a corresponding reverse-complement rule also exists within the de Bruijn set.
-The application of the properties stated in the previous paragraphs to the entire set of generating rules significantly reduces the number of feasible rules that can yield de Bruijn sequences.
+Let us summarize the reduction procedure that results from the application of these properties for the case $\mu=4$. There are 65,536 rules that are filtered as follows:
+  • After applying the boundary conditions, 16,384 rules remain feasible.
+  • Applying symmetry, i.e., dividing by $\phi(4)=255$, 64 rules remain.
+  • The requirement that the remainder is an Evil Odd Number reduces the previous number to 32 rules.
+  • Verifying the constrained position pairs reduces the feasible set to 24 rules.
+From these 24 feasible rules, only 16 are de Bruijn (see Table 2).
+The application of the properties stated in the previous paragraphs to the entire set of generating rules significantly reduces the number of feasible rules that can yield de Bruijn sequences (an example can be found in Appendix C).
 Table 3 presents the total number of rules, the number of de Bruijn sequences, the number of feasible rules after applying the constraints, and the corresponding ratios.
@@
 As $\mu$ increases, the reduction in the search space for de Bruijn rules becomes dramatic — for example, for $\mu=9$, the feasible subset constitutes only about $10^{-79}$ of the full rule set.
-$\mu$ $C(\mu)$#\## Feasible ##\## de Bruijn Feasible/Total de Bruijn/Total de Bruijn/Feasible
+$\mu$ $C(\mu)$ #\# Feasible #\# de Bruijn Feasible/Total de Bruijn/Total de Bruijn/Feasible
 2 16 1 0.0625
 3 256 2 2 0.0078125 0.0078125 1
-4 65536 24 16 0.000366211 0.000244141 0.66666667
-5 4294967296 6144 2048 1.4305E-06 4.7683E-07 0.33333333
-6 1.84467E+19 402653184 67108864 2.1827E-11 3.6379E-12 0.16666667
-7 3.40282E+38 1.7293E+18 1.4411E+17 5.082E-21 4.2351E-22 0.08333333
-8 1.15792E+77 3.1901E+37 1.3292E+36 2.7550E-40 1.1479E-41 0.04166667
-9 1.3408E+154 1.0855E+76 2.2615E+74 8.0964E-79 1.6867E-80 0.02083333
+4 65,536 24 16 0.000366211 0.00024414 0.66666667
+5 4,294,967,296 6144 2048 1.4305 $\times$ 106-6 4.7683 $\times$ 107-7 0.33333333
+6 1.84467 $\times$ 101919 402,653,184 67,108,864 2.1827 $\times$ 1011-11 3.6379 $\times$ 1012-12 0.16666667
+7 3.40282 $\times$ 103838 1.7293 $\times$ 101818 1.4411 $\times$ 101717 5.0820 $\times$ 1021-21 4.2351 $\times$ 1022-22 0.08333333
+8 1.15792 $\times$ 107777 3.1901 $\times$ 103737 1.3292 $\times$ 103636 2.7550 $\times$ 1040-40 1.1479 $\times$ 1041-41 0.04166667
+9 1.3408 $\times$ 10154154 1.0855 $\times$ 107676 2.2615 $\times$ 107474 8.0964 $\times$ 1079-79 1.6867 $\times$ 1080-80 0.02083333
 5 Neural Networks to Classify de Bruijn Rules
 An alternative approach to identifying de Bruijn rules is to apply machine learning methods.
-In particular, neural networks are especially well-suited for classification tasks Bishop .
+In particular, neural networks are especially well suited for classification tasks Bishop .
 In this section, we present a neural network model for classifying the feasible rules into two categories: de Bruijn rules (coded by 1) and the rest (0).
@@
 The classification is performed for $\mu=5$ and $\mu=6$ (see Table 4).
-We implemented a binary classification model using a feedforward neural network in R Rcore . The analysis followed these main steps:
+We implemented a binary classification model using a feedforward neural network in R (version 4.5.0) Rcore . The analysis followed these main steps:
   1. 1.
@@
   3. 3.
-    Dataset splitting: The data was randomly split into training (80%) and testing (20%) subsets to evaluate model performance on unseen data.
+    Dataset splitting: The data were randomly split into training (80%) and testing (20%) subsets to evaluate model performance on unseen data.
+The validation set was chosen to be $20\%$ of the training set in all cases.
   4. 4.
-    Model specification and training: A feedforward neural network was constructed using the keras package with a TensorFlow backend Allaire ; Rcore .
+    Model specification and training: A feedforward neural network was constructed using the keras package (version 2.15.0) with a TensorFlow (version 2.16.0) backend Allaire ; Rcore .
 Feedforward neural networks are highly suitable for structured, tabular data, such as the bit vectors extracted from rule representations.
@@
 The multi-layer architecture enables hierarchical feature learning, significantly improving classification accuracy.
-For the case $\mu=5$, the neural network architecture includes:
-  • An input layer with 14 features (bits),
-  • A hidden dense layer with 32 units and ReLU activation,
-  • A second hidden dense layer with 16 units and ReLU activation,
+For the case $\mu=5$, the neural network architecture includes the following:
+  • An input layer with 14 features (bits);
+  • A hidden dense layer with 32 units and ReLU activation;
+  • A second hidden dense layer with 16 units and ReLU activation;
   • An output layer with 1 unit and sigmoid activation for binary classification.
-The model was compiled using the Adam optimizer (learning rate $=0.001$), binary cross-entropy loss function, and accuracy as a performance metric.
+The model was compiled using the Adam optimizer (learning rate =0.001=0.001), binary cross-entropy loss function, and accuracy as a performance metric.
 The dataset includes the complete set of 6144 feasible rules, one-third of which are de Bruijn rules (see Table 3).
@@
 As a matter of fact, we can only use a sample of the total rule space that is randomly analyzed and classified into the two classes.
+To get this sample of the feasible rule space, we used brute force to generate 1 million de Bruijn rules.
+In parallel, we generated a larger number of non-de Bruijn feasible rules and randomly selected 1 million of them.
+Note that the ratio of de Bruijn to feasible rules is about 0.17 (Table 3), which means that approximately 6 million feasible rules must be generated by brute force to obtain 1 million de Bruijn rules.
 Optimal results were achieved using a slightly deeper neural network with the following architecture:
-  • Three hidden dense layers with 64, 64, and 8 units respectively,
-  • ReLU activations in all hidden layers,
+  • Three hidden dense layers with 64, 64, and 8 units, respectively;
+  • ReLU activations in all hidden layers;
   • A sigmoid output unit for binary classification.
@@
 The dataset consists of a sample of $2\times 10^{6}$ feasible rules, half of which are de Bruijn rules.
-As before, 80% of the dataset was used for training, and the remainder for testing.
+As before, 80% of the dataset was used for training, with 20% used for validation and the remainder for testing.
 A threshold of 0.5 was again applied to the output probabilities to assign class labels.
@@
 Table 4: Evaluation metrics for the neural network models applied to classify the de Bruijn rules for memories with $\mu=5$ and $\mu=6$.
-Total refers to the sum: $TP+FP+TN+FN$, and corresponds to one fith of the whole data set.
-Positive class considered: 1 $\to$ de Bruijn.
-Metric (definition) $\mu=5$ $\mu=6$
-True Positives (TP) 397 198563
-False Positives (FP) 3 19839
-True Negatives (TN) 820 180668
+Total refers to the sum TP+FP+TN+FN and corresponds to one fifth of the whole dataset.
+The positive class considered 1 $\to$ de Bruijn.
+Metric, Definition $\mu=5$ $\mu=6$
+True Positives (TP) 397 198,563
+False Positives (FP) 3 19,839
+True Negatives (TN) 820 180,668
 False Negatives (FN) 9 930
 Accuracy, (TP + TN)/Total 0.9902 0.9481
-Sensitivity (Recall), TP / (TP + FN) 0.9780 0.9953
+Sensitivity (Recall), TP/(TP + FN) 0.9778 0.9953
 Specificity, TN/(TN + FP) 0.9964 0.9011
@@
 Negative Predictive Value (NPV), TN/(TN + FN) 0.9891 0.9949
-Balanced Accuracy, (Sens. + Spec.) / 2 0.9872 0.9482
-Detection Rate, TP / Total 0.3230 0.4976
-Detection Prevalence, (TP + FP) / Total 0.3255 0.5473
-True Prevalence, (TP + FN) / Total 0.3304 0.4999
+Balanced Accuracy, (Sens. + Spec.)/2 0.9871 0.9482
+Detection Rate, TP/Total 0.3230 0.4964
+Detection Prevalence, (TP + FP) / Total 0.3255 0.5460
+True Prevalence, (TP + FN) / Total 0.3303 0.4987
 Once the neural network model is available, any given rule can be evaluated to determine, with high probability, whether it is a de Bruijn rule.
@@
 For large values of $\mu$, the number of possible generating rules becomes so large that an exhaustive analysis of the resulting patterns is infeasible.
-Particularly relevant is a specific and highly constrained subset of rules, that we named as de Bruijn rules, that generate sequences of maximum period, known as de Bruijn sequences.
+Particularly relevant is a specific and highly constrained subset of rules that we named as de Bruijn rules and that generate sequences of maximum period, known as de Bruijn sequences.
 Although these sequences represent only a tiny fraction of the entire set of possible sequences, their sheer number for large $\mu$ remains enormous, and the complete and effective generation of all such sequences is still an open problem.
 In this paper, we have presented a novel approach to compute de Bruijn sequences that combines two complementary methodologies.
-First, by exploiting structural properties of de Bruijn rules—i.e., those rules that generate maximum period sequences—a drastic reduction of the full set of candidate rules can be achieved.
+First, by exploiting structural properties of de Bruijn rules—i.e., those rules that generate maximum period sequences—a drastic reduction in the full set of candidate rules can be achieved.
 Table 3 summarizes the reduction ratios for several values of $\mu$.
 Second, we applied a machine learning approach to this feasible subset in order to accurately identify the de Bruijn rules.
-As shown in Section 5, the use of a classical neural network model for $\mu=5$, and $6$ allows for nearly complete classification of de Bruijn rules in the smaller cases, and achieves over 99% and 94% accuracy for $\mu=5$ and $\mu=6$, respectively.
+As shown in Section 5, the use of a classical neural network model for $\mu=5$ and $6$ allows for a nearly complete classification of the de Bruijn rules in the smaller cases and achieves over 99% and 94% accuracy for $\mu=5$ and $\mu=6$, respectively.
 Once the de Bruijn rules are identified, the corresponding de Bruijn sequences are straightforward to construct.
-This would allow to find granddaddy sequence that represents the lexicographically smallest member for each $\mu$-value.
-To the best of our knowledge, this is the first attempt to generate de Bruijn sequences using a hybrid approach that integrates Machine Learning techniques within the framework of Cellular Automata for generating maximum period sequences.
-It is worth mentioning that existing methods based on Feedback Shift Register generation have not definitively solved the problem.
-Therefore, we believe that the results presented in this paper represent a significant step forward in the study and generation of this important class of sequences.
+This would allow one to find the granddaddy sequence that represents the lexicographically smallest member for each $\mu$-value.
+The rule-based approach for generating de Bruijn sequences presented in this paper is, to the best of our knowledge, unique in the literature.
+It employs a hybrid methodology that integrates machine learning techniques within the framework of cellular automata.
+Specifically, we use rules with memory $\mu$ to generate sequences from an initial configuration of $\mu$ bits.
+Rules that generate sequences with a maximum period $T=2^{\mu}$—the so-called de Bruijn sequences—are referred to as the de Bruijn rules.
+By characterizing these rules based on certain empirical properties, we achieve a significant reduction in the overall rule set.
+However, even after this reduction, a huge set of feasible rules remains for larger $\mu$-values, which still needs to be classified.
+To address this, we employ a neural network, which has been shown to perform well for $\mu=5$ and $\mu=6$.
+We believe that further development of this machine learning methodology can yield excellent results for computing de Bruijn sequences with much larger $\mu$-values.
+It is also worth noting that existing methods based on Nonlinear Feedback Shift Registers have yet to definitively solve the problem.
+Thus, we consider the results presented in this paper a significant advancement in the study and generation of this important class of sequences.
+Appendix A
+In this appendix, we show that a natural number $x_{n}$ whose binary representation of 2n2\,n bits has the form
+$x_{n}\equiv[a_{1},a_{2},\dots,a_{n},1-a_{1},1-a_{2},\dots,1-a_{n}]$
+for $n\in\mathbb{N}$ and $a_{i}\in\{0,1\}$ are divisible by
+$M_{n}=2^{n}-1$
+The decimal representation of $x_{n}$ can be written as
+$x_{n}=\sum_{i=1}^{n}a_{i}\cdot 2^{2n-i}+\sum_{i=1}^{n}(1-a_{i})\cdot 2^{n-i}$
+which can be manipulated as follows:
+$x_{n}$ $=\sum_{i=1}^{n}\left(a_{i}\cdot 2^{2n-i}+(1-a_{i})\cdot 2^{n-i}\right)$
+$=\sum_{i=1}^{n}\left(a_{i}(2^{2n-i}-2^{n-i})+2^{n-i}\right)$
+$=\sum_{i=1}^{n}a_{i}\cdot 2^{n-i}(2^{n}-1)+\sum_{i=1}^{n}2^{n-i}$
+The last sum is geometric and can be summed as
+$\sum_{i=1}^{n}2^{n-i}=2^{n}-1.$
+and, denotating $A=\sum_{i=1}^{n}a_{i}\cdot 2^{n-i}$, the decimal expression of $x_{n}$ can be rewritten:
+$x_{n}=(2^{n}-1)(A+1)$
+Therefore, factor $M_{n}$ divides $x_{n}$ for $n\in\mathbb{N}$.
+This result can be applied to a de Bruijn rule with period $T=2^{\mu}$.
+In this case, $n=2^{\mu-1}$, and, then, $\phi(\mu)=M_{n}=2^{2^{\mu-1}}-1$.
+In this way, we obtain the factors that appear in Table 2:
+  • $n=4\Rightarrow\mu=3\quad\text{and}\quad\phi(3)=2^{4}-1=15=3\cdot 5$.
+  • $n=8\Rightarrow\mu=4\quad\text{and}\quad\phi(4)=2^{8}-1=255=3\cdot 5\cdot 17$.
+  • $n=16\Rightarrow\mu=5\quad\text{and}\quad\phi(5)=2^{16}-1=65535=3\cdot 5\cdot 17\cdot 257$.
+  • $n=32\Rightarrow\mu=6\quad\text{and}\quad\phi(6)=2^{32}-1=4294967295=3\cdot 5\cdot 17\cdot 257\cdot 65537$.
+The other factors $A+1$ are all Evil Odd Numbers A129771 as we have shown empirically for low values of $\mu$ (see Section 4).
+Therefore, the decimal representation of any de Bruijn rule for any $\mu$ can be decomposed into two terms, one being $\phi(\mu)$ and the other an Evil Odd Number.
+Appendix B
+The decimal representation of the de Bruijn rules for each $\mu$ is divisible for $\phi(\mu)$.
+The remainder of this division are Evil Odd Numbers.
+In Table 2, the Evil Odd Numbers corresponding to $\mu=2,$ 33, and $4$ are depicted.
+In this appendix, we show the Evil Odd Numbers that divide the de Bruijn rules for $\mu=5$.
+In total, there are 2048 Evil Odd Numbers, one for each de Bruijn rule.
+Table 5: This table continues Table 2 and depicts the subset of Evil Odd Numbers for $\mu=5$. The corresponding $\phi$-value is $\phi(5)=65,535$.
+39 2599 3751 6245 7565 9127 10,203 12,729 13,977 16,431 18,993 20,127 22,689 23,941 25,531 26,573 29,173 30,311
+43 2605 3763 6255 7571 9139 10,205 12,771 13,983 16,433 18,999 20,143 22,695 23,951 25,533 26,585 29,177 30,323
+45 2611 3789 6257 7573 9141 10,221 12,777 13,999 16,439 19,003 20,155 22,701 23,963 25,581 26,659 29,223 30,341
+51 2617 3801 6263 7589 9199 10,233 12,785 14,011 16,551 19,053 20,175 22,713 23,965 25,593 26,665 29,229 30,353
+57 2623 3807 6267 7593 9211 10,403 12,837 14,021 16,555 19,065 20,187 22,719 23,969 25,601 26,683 29,235 30,359
+63 2671 3823 6309 7599 9213 10,409 12,849 14,031 16,557 19,071 20,231 22,755 23,975 25,607 26,727 29,241 30,375
+165 2683 3835 6313 7601 9219 10,427 12,855 14,033 16,563 19,111 20,237 22,757 23,981 25,613 26,731 29,247 30,387
+169 2725 3845 6319 7611 9225 10,471 12,903 14,039 16,569 19,117 20,243 22,767 23,987 25,625 26,733 29,285 30,407
+175 2735 3855 6321 7613 9243 10,475 12,915 14,043 16,575 19,123 20,245 22,769 23,989 25,631 26,739 29,295 30,419
+177 2737 3857 6327 7619 9255 10,477 12,967 14,055 16,679 19,129 20,249 22,775 23,993 25,635 26,745 29,297 30,469
+183 2743 3867 6369 7625 9259 10,483 12,973 14,067 16,683 19,135 20,261 22,779 24,001 25,637 26,751 29,303 30,479
+293 2747 3869 6379 7633 9261 10,489 12,979 14,087 16,685 19,183 20,273 22,823 24,007 25,647 26,913 29,307 30,481
+297 2797 3885 6387 7649 9267 10,495 12,985 14,099 16,691 19,195 20,303 22,827 24,013 25,649 26,923 29,349 30,491
+303 2809 3897 6435 7655 9273 10,657 12,991 14,101 16,693 19,239 20,315 22,829 24,019 25,655 26,937 29,361 30,493
+305 2815 3917 6437 7661 9279 10,667 13,029 14,117 16,805 19,245 20,317 22,835 24,021 25,659 26,981 29,367 30,509
+315 2853 3929 6447 7667 9287 10,681 13,039 14,129 16,809 19,251 20,333 22,837 24,025 25,669 26,985 29,415 30,521
+317 2863 3975 6459 7669 9291 10,725 13,041 14,149 16,815 19,253 20,345 22,883 24,077 25,673 26,991 29,427 30,535
+423 2865 3981 6461 7673 9293 10,729 13,047 14,161 16,817 19,257 20,357 22,889 24,089 25,679 26,993 29,477 30,541
+427 2875 3987 6497 7685 9299 10,735 13,051 14,213 16,827 19,311 20,367 22,897 24,095 25,681 27,003 29,487 30,547
+429 2877 3989 6503 7697 9305 10,737 13,095 14,223 16,829 19,323 20,369 22,947 24,111 25,687 27,005 29,489 30,549
+435 2925 3993 6509 7703 9311 10,747 13,107 14,225 16,941 19,325 20,379 22,949 24,123 25,703 27,175 29,499 30,553
+437 2937 4005 6515 7719 9345 10,749 13,109 14,235 16,953 19,365 20,381 22,959 24,133 25,707 27,181 29,501 30,565
+559 2983 4017 6517 7731 9351 10,919 13,157 14,237 16,959 19,375 20,397 22,971 24,143 25,709 27,187 29,543 30,577
+571 2989 4047 6521 7751 9357 10,925 13,169 14,253 17,071 19,377 20,409 22,973 24,145 25,715 27,193 29,549 30,599
+685 2995 4059 6567 7763 9369 10,931 13,221 14,265 17,083 19,387 20,429 23,009 24,151 25,721 27,199 29,555 30,611
+697 2997 4061 6571 7821 9375 10,937 13,231 14,279 17,199 19,389 20,441 23,015 24,155 25,727 27,247 29,557 30,613
+703 3001 4077 6573 7833 9379 10,943 13,233 14,285 17,211 19,437 20,513 23,021 24,167 25,731 27,259 29,561 30,629
+813 3055 4089 6579 7839 9381 10,991 13,243 14,291 17,213 19,449 20,523 23,027 24,179 25,737 27,429 29,607 30,641
+825 3067 4131 6581 7855 9391 11,003 13,245 14,293 17,325 19,457 20,531 23,029 24,197 25,755 27,439 29,619 30,661
+943 3069 4133 6627 7867 $93$ 11,173 13,287 14,297 17,337 19,467 20,643 23,033 24,209 25,767 27,441 29,621 30,673
+955 3075 4143 6633 7877 9399 11,183 13,293 14,309 17,413 19,475 20,645 23,085 24,215 25,771 27,451 29,669 30,753
+957 3081 4145 6641 7887 9403 11,185 13,299 14,321 17,417 19,491 20,655 23,097 24,231 25,773 27,453 29,681 30,759
+1031 3099 4151 6693 7889 9413 11,195 13,301 14,497 17,423 19,493 20,657 23,103 24,243 25,779 27,501 29,699 30,765
+1035 3111 4155 6705 7895 9417 11,197 13,305 14,503 17,425 19,503 20,663 23,141 24,263 25,785 27,513 29,705 30,777
+1037 3115 4257 6711 7899 9423 11,245 13,313 14,509 17,431 19,505 20,667 23,151 24,275 25,791 27,651 29,723 30,783
+1043 3117 4267 6759 7911 9425 11,257 13,319 14,521 17,447 19,511 20,771 23,153 24,335 25,799 27,657 29,735 30,819
+1049 3123 4275 6771 7923 9431 11,395 13,325 14,527 17,451 19,515 20,777 23,159 24,347 25,803 27,675 29,739 30,821
+1055 3129 4385 6829 7943 9447 11,401 13,337 14,563 17,453 19,525 20,785 23,163 24,349 25,805 27,687 29,741 30,831
+1157 3135 4391 6841 7955 9451 11,419 13,343 14,565 17,459 19,529 20,897 23,205 24,365 25,811 27,691 29,747 30,833
+1161 3143 4397 6847 7957 9453 11,431 13,347 14,575 17,465 19,535 20,903 23,217 24,377 25,817 27,693 29,753 30,839
+1167 3147 4403 6885 7973 9459 11,435 13,349 14,577 17,471 19,537 20,909 23,223 24,391 25,823 27,699 29,759 30,843
+1169 3149 4405 6895 7985 9465 11,437 13,359 14,583 17,543 19,543 20,915 23,271 24,397 25,859 27,705 29,761 31,011
+1175 3155 4409 6897 8005 9471 11,443 13,361 14,587 17,547 19,559 20,917 23,283 24,403 25,861 27,711 29,771 31,013
+1191 3161 4515 6903 8017 9473 11,449 13,367 14,755 17,549 19,563 20,921 23,343 24,405 25,871 27,719 29,779 31,023
+1195 3167 4521 6907 8079 9483 11,455 13,371 14,757 17,555 19,565 21,029 23,355 24,409 25,883 27,723 29,795 31,035
+1197 3201 4529 6951 8091 9497 11,463 13,379 14,767 17,561 19,571 21,039 23,357 24,421 25,885 27,725 29,797 31,037
+1203 3211 4647 6963 8093 9509 11,467 13,381 14,779 17,567 19,577 21,041 23,399 24,433 25,889 27,731 29,807 31,073
+1209 3219 4659 6965 8109 9513 11,469 13,391 14,781 17,671 19,583 21,047 23,405 24,455 25,895 27,737 29,809 31,079
+1215 3235 4773 7013 8121 9519 11,475 13,393 14,817 17,675 19,587 21,051 23,411 24,467 25,901 27,743 29,815 31,085
+1285 3237 4783 7025 8135 9521 11,481 13,399 14,823 17,677 19,593 21,159 23,413 24,469 25,907 27,905 29,819 31,091
+1289 3247 4785 7087 8141 9531 11,487 13,403 14,829 17,683 19,611 21,171 23,417 24,485 25,909 27,915 29,825 31,093
+1295 3249 4791 7099 8147 9533 11,649 13,443 14,835 17,685 19,623 21,287 23,463 24,497 25,913 27,929 29,831 31,097
+1297 3255 4795 7101 8149 9541 11,659 13,449 14,837 17,701 19,627 21,293 23,475 24,517 25,927 27,941 29,837 31,269
+1307 3259 4901 7143 8153 9545 11,673 13,467 14,841 17,705 19,629 21,299 23,477 24,529 25,931 27,945 29,849 31,281
+1309 3269 4913 7149 8165 9551 11,685 13,479 15,013 17,711 19,635 21,301 23,525 24,609 25,933 27,951 29,855 31,287
+1415 3273 5031 7155 8177 9553 11,689 13,483 15,025 17,713 19,641 21,305 23,537 24,615 25,939 27,953 29,859 31,335
+1419 3279 5037 7157 8227 9563 11,695 13,485 15,031 17,723 19,647 21,413 23,557 24,621 25,941 27,963 29,861 31,347
+1421 3281 5043 7161 8233 9565 11,697 13,491 15,079 17,725 19,655 21,425 23,561 24,633 25,957 27,965 29,871 31,527
+1427 3287 5045 7169 8251 9603 11,707 13,497 15,091 17,797 19,659 21,505 23,567 24,639 25,961 27,973 29,873 31,539
+1429 3303 5049 7175 8295 9605 11,709 13,503 15,271 17,801 19,661 21,515 23,569 24,677 25,967 27,977 29,879 31,541
+1445 3307 5123 7181 8299 9615 11,717 13,505 15,283 17,807 19,667 21,523 23,575 24,681 25,969 27,983 29,883 31,589
+1449 3309 5125 7193 8301 9627 11,721 13,515 15,285 17,809 19,673 21,539 23,591 24,687 25,979 27,985 29,891 31,601
+1455 3315 5135 7199 8307 9629 11,727 13,523 15,333 17,819 19,679 21,541 23,595 24,689 25,981 27,995 29,893 31,745
+1457 3321 5137 7203 8313 9633 11,729 13,539 15,345 17,821 19,715 21,551 23,597 24,695 25,985 27,997 29,903 31,751
+1467 3327 5143 7205 8319 9639 11,739 13,541 15,489 17,933 19,721 21,553 23,603 24,739 25,995 28,167 29,905 31,757
+1469 3329 5147 7215 8353 9645 11,741 13,551 15,495 17,945 19,729 21,559 23,609 24,745 26,009 28,173 29,911 31,769
+1551 3339 5249 7217 8359 9651 11,911 13,553 15,501 17,951 19,745 21,563 23,615 24,763 26,021 28,179 29,915 31,775
+1563 3353 5259 7223 8365 9653 11,917 13,559 15,513 17,967 19,751 21,635 23,617 24,807 26,025 28,185 29,953 31,779
+1677 3365 5267 7227 8377 9657 11,923 13,563 15,519 17,979 19,757 21,637 23,627 24,811 26,031 28,191 29,963 31,781
+1689 3369 5283 7235 $83$ 9671 11,929 13,571 15,523 18,063 19,763 21,647 23,635 24,813 26,033 28,207 29,977 31,791
+1695 3375 5285 7237 8421 9675 11,935 13,573 15,525 18,075 19,765 21,649 23,651 24,819 26,043 28,219 29,989 31,793
+1711 3377 5295 7247 8425 9677 11,951 13,583 15,535 18,191 19,769 21,655 23,653 24,825 26,045 28,239 29,993 31,799
+1723 3387 5297 7249 8431 9683 11,963 13,595 15,537 18,203 19,783 21,659 23,663 24,831 26,053 28,251 29,999 31,803
+1805 3389 5303 7255 8433 9685 11,983 13,597 15,543 18,205 19,787 21,763 23,665 24,867 26,057 28,421 30,001 31,811
+1817 3397 5307 7259 8439 9701 11,995 13,601 15,547 18,221 19,789 21,769 23,671 24,869 26,063 28,431 30,011 31,813
+1935 3401 5377 7301 8481 9705 12,165 13,607 15,555 18,233 19,795 21,777 23,675 24,879 26,065 28,433 30,013 31,823
+1947 3407 5383 7305 8491 9711 12,175 13,613 15,557 18,317 19,797 21,793 23,681 24,891 26,075 28,443 30,019 31,825
+1949 3409 5389 7311 8505 9713 12,177 13,619 15,567 18,329 19,813 21,799 23,687 24,893 26,077 28,445 30,025 31,831
+1965 3419 5395 7313 8549 9723 12,187 13,621 15,569 18,465 19,817 21,805 23,693 24,935 26,117 28,461 30,033 31,835
+1977 3421 5397 7319 8553 9725 12,189 13,625 15,575 18,475 19,823 21,811 23,705 24,939 26,129 28,473 30,049 32,003
+2083 3459 5401 7335 8559 9735 12,205 13,633 15,579 18,483 19,825 21,813 23,711 24,941 26,135 28,493 30,055 32,005
+Table 5: Cont.
+2089 3465 5507 7339 8561 9741 12,217 13,639 15,747 18,533 19,835 21,817 23,715 24,947 26,151 28,505 30,061 32,015
+2107 3473 5513 7341 8571 9747 12,237 13,645 15,749 18,537 19,837 21,889 23,717 24,949 26,163 28,707 30,067 32,027
+2151 3489 5521 7347 8573 9753 12,249 13,651 15,759 18,543 19,841 21,895 23,727 24,993 26,189 28,713 30,069 32,029
+2155 3495 5537 7353 8611 9759 12,321 13,653 15,771 18,545 19,851 21,901 23,729 25,003 26,201 28,731 30,073 32,033
+2157 3501 5543 7359 8613 9775 12,327 13,657 15,773 18,551 19,865 21,907 23,735 25,017 26,207 28,769 30,083 32,039
+2163 3507 5549 7361 8623 9787 12,333 13,697 15,777 18,595 19,877 21,909 23,739 25,061 26,223 28,779 30,085 32,045
+2169 3509 $55$ 7371 8635 9807 12,345 13,707 15,783 18,601 19,881 21,913 23,747 25,065 26,235 28,787 30,095 32,051
+2175 3513 5557 7379 8637 9819 12,351 13,721 15,789 18,619 19,887 22,021 23,749 25,071 26,247 28,833 30,107 32,053
+2209 3527 5561 7395 8679 9861 12,387 13,733 15,795 18,663 19,889 22,031 23,759 25,073 26,253 28,839 30,109 32,057
+2219 3531 5639 7397 8683 9873 12,389 13,737 15,797 18,667 19,899 22,033 23,761 25,083 26,259 28,845 30,113 32,065
+2227 3533 5651 7407 8685 9879 12,399 13,743 15,801 18,669 19,901 22,039 23,767 25,085 26,265 28,857 30,119 32,071
+2277 3539 5765 7409 8691 9895 12,401 13,745 15,809 18,675 19,909 22,043 23,771 25,125 26,271 28,863 30,125 32,077
+2281 3541 5775 7415 8693 9907 12,407 13,755 15,815 18,681 19,913 22,055 23,815 25,137 26,287 28,899 30,131 32,083
+2287 3557 5777 7419 8743 9933 12,411 13,757 15,821 18,687 19,919 22,067 23,819 25,143 26,299 28,901 30,133 32,085
+2289 3561 5783 7427 8749 9945 12,451 13,763 15,827 18,723 19,921 22,151 23,821 25,197 26,319 28,911 30,137 32,089
+2295 3567 5787 7429 8755 9951 12,457 13,769 15,829 18,729 19,931 22,163 23,827 25,209 26,331 28,913 30,145 32,261
+2337 3569 5799 7439 8761 9967 12,475 13,777 15,833 18,737 19,933 22,279 23,829 25,215 26,375 28,919 30,151 32,273
+2347 3579 5811 7451 8767 9979 12,513 13,793 16,005 18,791 19,973 22,285 23,845 25,255 26,387 28,923 30,157 32,279
+2361 3581 5893 7453 8815 9989 12,523 13,799 16,017 18,795 19,983 22,291 23,849 25,261 26,389 28,961 30,163 32,295
+2405 3591 5905 7457 8827 $99$ 12,531 13,805 16,023 18,797 19,985 22,293 23,855 25,267 26,405 28,971 30,165 32,307
+2409 3597 6023 7463 8869 10,001 12,579 13,811 16,039 18,803 19,991 22,297 23,857 25,273 26,417 28,985 30,169 32,327
+2415 3603 6029 7469 8881 10,011 12,581 13,813 16,051 18,805 19,995 22,309 23,867 25,279 26,447 29,027 30,215 32,339
+2417 3609 6035 7475 8887 10,013 12,591 13,817 16,071 18,849 20,007 22,321 23,869 25,327 26,459 29,033 30,221 32,519
+2427 3615 6037 7477 8941 10,029 12,603 13,829 16,083 18,859 20,019 22,405 23,875 25,339 26,461 29,041 30,227 32,531
+2429 3631 6041 7481 8953 10,041 12,605 13,841 16,263 18,873 20,045 22,417 23,881 25,383 26,477 29,091 30,233 32,533
+2467 3643 6053 7489 8959 10,061 12,641 13,847 16,275 18,917 20,057 22,565 23,889 25,395 26,489 29,093 30,239 32,549
+2473 3663 6065 7495 8997 10,073 12,647 13,863 16,277 18,921 20,063 22,569 23,905 25,397 26,501 29,103 30,255 32,561
+2481 3675 6177 7501 9007 10,119 12,653 13,875 16,293 18,927 20,079 22,575 23,911 25,455 26,511 29,115 30,267 32,581
+2535 3717 6183 7507 9009 10,131 12,659 13,895 16,305 18,929 20,091 22,577 23,917 25,467 26,513 29,117 30,277 32,593
+2539 3727 6189 7509 9019 10,133 12,661 13,907 16,325 18,939 20,103 22,583 23,923 25,469 26,523 29,153 30,287
+2541 3729 6201 7513 9021 10,149 12,665 13,959 16,337 18,941 20,109 22,625 23,925 25,509 26,525 29,159 30,289
+2547 3735 6207 7559 9069 10,161 12,705 13,965 16,421 18,981 20,115 22,635 23,929 25,519 26,541 29,165 30,295
+2549 3739 6243 7563 9081 10,191 12,715 13,971 16,425 18,991 20,121 22,643 23,939 25,521 26,553 29,171 30,299
+Appendix C
+As an application of the methodology presented in this paper, we can generate sequences with large periods, $T=2^{\mu}$, by tuning the memory parameter $\mu$.
+We have proven that, once $\mu$ is chosen, all de Bruijn rules share the same divisor: $\phi(\mu)=2^{2^{\mu-1}}-1$, and the remainders are Evil Odd Numbers.
+Therefore, to obtain a candidate to de Bruijn rule, take a very large Evil Odd Number and multiply it by $\phi(\mu)$.
+Check the resulting rule and, if valid, apply the rule to any initial binary sequence of length $\mu$ to obtain the de Bruijn sequence.
+Due to computational limitations, we can only offer an example for $\mu=8$, which provides a sequence of period of $T=256$. For this case,
+$\phi(8)=2^{2^{7}}-1=340282366920938463463374607431768211455$
+and take the Evil Odd Number 155640396308704138405661716133499575.
+Multiply both to obtain the decimal representation of the de Bruijn rule:
+$R_{8}=52961682444438738040038706846647584952275525588128484521019054507752631625$
+The binary representation of this de Bruijn rule is
+0000000000011101111110011010100111111011100101001010000100111111
+1001100000011011011001001111010110000001000001001101001010110110
+1111111111100010000001100101011000000100011010110101111011000000
+0110011111100100100110110000101001111110111110110010110101001001
+Applying this rule to the initial configuration 00000000 yields the following de Bruijn sequence:
+0000000010001111101100111000111011011100100010000110000001110011
+0001011010101011011000110011011010011010000010101110100010100111
+1000100111010111100111111010010111000001101111010100100110010100
+0110101100001111011101111111100001011111001011001001010100001001
+This sequence of period $T=256$ repeats infinitely under the application of the above de Bruijn rule.